# FedFACT: A Provable Framework for Controllable Group-Fairness Calibration in Federated Learning

### v1

@@ B.8.1 Proof of Theorem 10
which can be regarded as the intersection of $m-1$ half-spaces given $\eta_{k,i},a_{k,i}$. A single halfspace function class can be viewed as the class of linear classifiers, possessing a VC dimension of $m$. By the additive property of VC dimension, for function classes $\{\mathcal{G}_{i}\}_{i=1}^{m}$, $VC\left(\bigwedge_{i=1}^{m}\mathcal{G}_{i}\right)\leq\sum_{i=1}^{m}VC\left(\mathcal{G}_{i}\right)$, the function class $\mathcal{H}_{i,j,a^{\prime}}$ has VC dimension at most $\mathcal{O}(\lvert\mathcal{A}\rvert m^{2})$. By taking a union bound in the Lemma B.3, we have that with probability at least $1-\delta$,
@@
${\mathscr{D}}^{g}_{u_{g}}({\widehat{\mathbf{h}}})-\widehat{\mathscr{D}}^{g}_{u_{g}}(\widehat{\mathbf{h}})$ $\leq\mathcal{O}(\rho(n,\delta/|\mathcal{A}|))|\mathcal{A}|NB_{g}+\Omega^{g}_{n}.$
@@
$\frac{\partial}{\partial\lambda^{(1)}_{u_{g}}}\widehat{H}(\lambda,\mu)=$ $\sum_{k\in[N]}\sum_{a\in\mathcal{A}}\hat{p}_{a,k}\frac{1}{n_{a,k}}\sum_{i=1}^{n_{a,k}}\Big{[}\frac{\partial}{\partial\lambda^{(1)}_{u_{g}}}\max_{y\in[m]}\Big{(}\big{[}\widehat{\mathbf{M}}^{\lambda,\mu}(a,k)\big{]}^{\top}\widehat{\eta}(x_{i},a,k)\Big{)}_{y}\Big{]}+\xi^{g}$ (55)
@@
Hence, the subgradient falls into an interval. Since $[\widehat{\eta}(x_{i},a,k)]^{\top}\widehat{\mathbf{D}}^{a,k}_{u_{g}}e_{t}\leq\|[\widehat{\eta}(x_{i},a,k)]^{\top}\widehat{\mathbf{D}}^{a,k}_{u_{g}}\|_{1}\leq B_{g}$, denoting $n_{\min}:=\min_{a\in\mathcal{A},k\in[N]}n_{a,k}$ and $\widehat{B}_{g}$, we have that
@@
Since we assume that $n>\frac{2|\mathcal{A}|N\widehat{B}_{l,k}}{p_{\min}\xi^{g}}+\frac{1}{2p_{\min}^{2}}\log\frac{1}{\delta}$, by Lemma B.7, we have that with probability at last $1-|\mathcal{A}|\delta$,
@@
For local fairness constraints ${\mathscr{D}}^{l,k}({\widehat{\mathbf{h}}^{*}})$, following the same proof procedures, we arrive at that with probability at least $1-2|\mathcal{A}|\delta$,
$|{\mathscr{D}}^{l,k}({\widehat{\mathbf{h}}^{*}})|$ $\leq\xi^{l,k}+\mathcal{O}(\rho(n,N\delta/|\mathcal{A}||\mathcal{U}_{g}|))|\mathcal{A}|B_{l,k}+\Omega^{l,k}_{n}+\frac{|\mathcal{A}|\widehat{B}_{l,k}}{n_{\min}}.$
@@ C.1.1 Datasets
  • The Compas dataset [20] comprises 6,172 criminal defendants from Broward County, Florida, between 2013 and 2014, with the task of predicting whether a defendant will recidivate within two years of their initial risk assessment. We consider the race of each individual as the sensitive attribute and train a logistic classifier as our prediction model.
  • The Adult dataset [4] comprises more than 45000 samples based on 1994 U.S. census data, where the task is to predict whether the annual income of an individual is above \$50,000. We consider the gender of each individual as the sensitive attribute and train the logistic regression as the classification model.
  • The ENEM dataset [36] contains about 1.4 million samples from Brazilian college entrance exam scores along with student demographic information. We follow [3] to quantized the exam score into 2 or 5 classes as label, and consider race as sensitive attribute. As [3] used a random subset of 50K samples, we instead sample 100K data points to construct our federated dataset. We train multilayer perceptron (MLP) as the classification model.
  • The CelebA dataset [83] is a facial image dataset consists of about 200k instances with 40 binary attribute annotations. We identify the binary feature smile as target attributes which aims to predict whether the individuals in the images exhibit a smiling expression. The race of individuals is chosen as sensitive attribute. We train Resnet18 [34] on CelebA as the classification model.
The determination of sensitive attributes and labels on three datasets has been verified significant in previous research [3, 31].
@@ C.1.2 Baselines
We compare the performance of FedFACT with traditional FedAvg [50] and five SOTA methods tailored for calibrating global and local fairness in FL, namely FairFed [29], FedFB [81], FCFL [17], praFFL [77], and the method in [23], denoted as Cost in our experiments.
@@
  • FedFB presents a FairBatch-based approach [60] to compute the coefficients of FairBatch parameters on the server. This method integrates global reweighting for each client into the FedAvg framework to fulfill fairness objectives.
@@
  • [23] proposed a convex-programming-based post-processing framework that characterizes and enforces the minimum accuracy loss required to satisfy specified levels of both local and global fairness constraints in multi-class federated learning by approximating the region under the ROC hypersurface with a simplex and solving a linear program, denoted as Cost in our experiments.
@@ C.1.3 Parameter Settings
For the fairness-control parameters, e.g., the parameter $\lambda$ in praFFL [77] and the global and local fairness constraints in Cost [23], we impose stringent fairness requirements on the model in our overall comparative experiments, and we adjust the parameters governing the fairness metrics in the Pareto-curve experiments.
@@ C.1.4 Experiments Compute Resources
C.2 Discussion about FedFACT and LoGoFair [82]
LogoFair [82] is designed for binary-classification in federated learning under both global and local fairness constraints, seeking the Bayes-optimal classifier. By deriving a closed-form solution for the fair Bayes classifier, LogoFair reformulates the post-processing fairness adjustment as a bilevel optimization problem jointly solved by the server and clients, which is an approach conceptually analogous to our post-processing framework. In binary classification, FedFact and LogoFair both target Bayes-optimal classifiers under constraints disparity metrics expressed in linear form. Theoretically, for an identical fairness metric, our Bayes-optimal fair classifier characterization covers that of LogoFair. Consequently, we refrain from performing a comparative evaluation of the two approaches.
Our method differs by defining the loss at the client level, thereby achieving lower estimation error than the local group-specific objective in [82]. Crucially, by formulating the post-processing model over the probabilistic simplex instead of restricting outputs to the unit interval $[0,1]$ in the binary case, our framework achieves enhanced scalability and naturally adaptable to multi-group, multiclass settings.
@@ C.3 Heterogeneous Split of Client Distribution
We propose a partitioning method that introduces heterogeneous correlations between the sensitive attribute $A$ and label $Y$, thereby further elucidating the trade-off between global fairness and local fairness [30].
@@
[…]
@@ Appendix D Detailed Experiments Results
D.1 Comparison Result and binary EO criterion
Parato Curves of DP. We have already presented the numerical comparison between our proposed method and the baselines in the main text; here, we report the Pareto curves illustrating the trade-off between global fairness and accuracy. More precisely, we compare the trade-off between accuracy and the global fairness measure, as well as the trade-off between accuracy and the local fairness measure, as a function of the fairness constraint.
The Pareto curve for the global DP criterion is shown in Figure 1 where the horizontal axis denotes accuracy and the vertical axis represents the fairness metric. Consequently, models located closer to the upper-right corner exhibit superior accuracy-fairness trade-offs. As illustrated in Figure 1, our method outperforms all existing state-of-the-art approaches when comparing accuracy against either global fairness in isolation.
@@ D.1 Comparison Result and binary EO criterion
Parato Curves of EO. In Figure 2, we illustrate the Pareto curve for the Equalized Odds (EO) criterion-accuracy. Because EO enforces tighter constraints than DP, precise adherence in a federated context requires large per-group sample counts at each client. Hence, we also compare the global EO here. Our framework still exceeds all state-of-the-art baselines in trading off accuracy against fairness.
@@
D.2 Result for multi-class classification
Multi-Class fair datasets. We illustrate how FedFACT performs on multi-class prediction using CelebA and ENEM. For CelebA, with ’Gender’ still serving as the sensitive attribute, We employ the binary attributes “Smile” and “Big_Nose” to construct a multiclass task by mapping their joint values $\{0,1\}\times\{0,1\}$ onto a four-class label set $\{0,1,2,3\}$, thereby formulating a multiclass classification problem on the CelebA dataset. These attributes are commonly used in centralized machine learning literature [12, 84] to construct fairness-aware classification tasks. For ENEM, we follow [3] to quantize the Humanities exam score to 5 classes. In order to guarantee adequate per-group sample sizes at each client in heterogeneous settings for fairness evaluation (or some clients only hold less than 10 samples for specific group under heterogeneous partitioning), we adopt the four race labels “Branca,” “Preta,” “Parda,” and “Amarela” from the Race attribute as the sensitive groups. These datasets are partitioned into five clients under a heterogeneous split with $\gamma=1$.
Evaluation. In terms of baselines, only the Cost [23] algorithm is theoretically applicable to fairness optimization in multiclass federated learning scenarios. However, their experiments and code are limited to binary classification, and have already been used as baselines for comparison with our method. Consequently, we focus exclusively on reporting FedFACT’s performance along with FedAvg in multiclass fairness, establishing it as a pioneering approach in this setting. In Figure 3, we plot the global-local fairness-accuracy trade-off of FedFACT in the multi-class classification task.
Figure 3: Multi-Class Fair Classification Results. Multiclass fairness calibration experiments in federated learning on the CelebA and ENEM datasets. The top panels depict global and local multiclass Demographic Parity (DP) results, while the bottom panels show global and local multiclass Equal Opportunity (EOP) outcomes.
Multi-Class Results. The outcomes of the multiclass experiments do not parallel those in the binary setting, where post-processing methods vastly outperform alternatives; instead, performance is comparatively lower. This can be attributed to the paucity of local samples for fairness evaluation at individual clients, which causes post-processing under joint global and local fairness constraints to incur significant generalization error and thus fail to precisely enforce local fairness (e.g. multiclass local EOP experiments in Figure 3). Under these conditions, the in-training approach, leveraging globally aggregated data, offers superior fairness calibration, thereby underscoring the complementarity of the two methods we introduce.
@@ D.5 Efficiency and Scalability Study
Efficiency. We evaluate the communication efficiency of FedFACT by monitoring its performance across varying numbers of communication rounds $T$. As illustrated in Figure 4, the post-processing method, built upon a fully trained pre-trained model, consistently achieves convergence in fewer than 10 communication rounds, underscoring its high efficiency. The in-processing method likewise converges in under 40 iterations; given that it requires training the federated model from scratch, this performance is comparable to the convergence speed of FedAvg, making it highly effective compared to existing federated learning algorithms.
Figure 4: Communication Effectiveness Analysis. The convergence rates of both the in-processing (top row) and post-processing (bottom row) methods with respect to communication rounds on Compas, Adult, CelebA, and ENEM datasets.
@@
Overall, whether employing the in-processing or post-processing method, all three performance metrics rapidly converge to stable values across each of the four datasets, empirically confirming both the communication efficiency and the overall effectiveness of FedFACT.
@@ Appendix E Broader Impacts and Limitations
Broader Impacts. This paper addresses critical fairness issues in FL. By embedding fairness constraints at both the global and client levels, our framework delivers models that distribute accuracy more equitably, bolstering user confidence and mitigating bias amplification. The contributions of this research enhance user satisfaction and promote social equity. This fairness-aware approach extends readily to high-stakes classification tasks beyond FL: for instance, clinical decision support in hospital networks, vision-based detection systems, and financial fraud alerts. Integrating fairness into decentralized model training promotes privacy-preserving, equitable AI, helps satisfy emerging regulatory requirements, and encourages broader adoption of responsible machine learning across diverse application domains.
Limitations. The primary limitation of FedFACT is the fairness representation, which contains the linear disparities such as communly used DP, EOP and EO criteria, but it excludes some nonlinear formulations of fairness, e.g. Predictive Parity [20] and individual fairness [27]. Moreover, based on our generalization-error analysis, although the proposed method enables a controllable accuracy-fairness trade-off for a given fairness metric, it still requires a sufficiently large local sample size to accurately estimate local fairness (whereas global fairness demands only an adequate overall sample size). While our empirical results compare favorably against existing approaches, exploiting dataset characteristics to optimize fairness may reduce the sample complexity needed for local fairness optimization. Addressing these limitations remains an important avenue for future work.

### v2

@@ B.8.1 Proof of Theorem 10
which can be regarded as the intersection of m1m-1 half-spaces given $\eta_{k,i},a_{k,i}$. A single halfspace function class can be viewed as the class of linear classifiers, possessing a VC dimension of $m$. By the additive property of VC dimension, for function classes $\{\mathcal{G}_{i}\}_{i=1}^{m}$, $VC\left(\bigwedge_{i=1}^{m}\mathcal{G}_{i}\right)\leq\sum_{i=1}^{m}VC\left(\mathcal{G}_{i}\right)$, the function class $\mathcal{H}_{i,j,a^{\prime}}$ has VC dimension at most $\mathcal{O}(\lvert\mathcal{A}\rvert m^{2})$. By taking a union bound in the Lemma B.3, we have that with probability at least $1-\delta$,
@@
[…]
@@
$\frac{\partial}{\partial\lambda^{(1)}_{u_{g}}}\widehat{H}(\lambda,\mu)=$ $\sum_{k\in[N]}\sum_{a\in\mathcal{A}}\hat{p}_{a,k}\frac{1}{n_{a,k}}\sum_{i=1}^{n_{a,k}}\Big[\frac{\partial}{\partial\lambda^{(1)}_{u_{g}}}\max_{y\in[m]}\Big(\big[\widehat{\mathbf{M}}^{\lambda,\mu}(a,k)\big]^{\top}\widehat{\eta}(x_{i},a,k)\Big)_{y}\Big]+\xi^{g}.$ (55)
@@
Hence, the subgradient falls into an interval. Since $\widehat{\eta}(x_{i},a,k)]^{\top}\widehat{\mathbf{D}}^{a,k}_{u_{g}}e_{t}\leq\|[\widehat{\eta}(x_{i},a,k)]^{\top}\widehat{\mathbf{D}}^{a,k}_{u_{g}}\|_{1}\leq B_{g}$, denoting $n_{\min}:=\min_{a\in\mathcal{A},k\in[N]}n_{a,k}$ and $\widehat{B}_{g}$, we have that
@@
Since we assume that $n>\frac{2|\mathcal{A}|N\widehat{B}_{k}}{p_{\min}\xi^{g}}+\frac{1}{2p_{\min}^{2}}\log\frac{1}{\delta}$, by Lemma B.7, we have that with probability at last $1-|\mathcal{A}|\delta$,
@@
For local fairness constraints ${\mathscr{D}}^{k}({\widehat{\mathbf{h}}^{*}})$, following the same proof procedures, we arrive at that with probability at least $1-2|\mathcal{A}|\delta$,
$|{\mathscr{D}}^{k}({\widehat{\mathbf{h}}^{*}})|$ $\leq\xi^{k}+\mathcal{O}(\rho(n,N\delta/|\mathcal{A}||\mathcal{U}_{g}|))|\mathcal{A}|B_{k}+\Omega^{k}_{n}+\frac{|\mathcal{A}|\widehat{B}_{k}}{n_{\min}}.$
@@ C.1.1 Datasets
  • The Compas dataset [compas-dressel2018accuracy] comprises 6,172 criminal defendants from Broward County, Florida, between 2013 and 2014, with the task of predicting whether a defendant will recidivate within two years of their initial risk assessment. We consider the race of each individual as the sensitive attribute and train a logistic classifier as our prediction model.
  • The Adult dataset [adult-asuncion2007uci] comprises more than 45000 samples based on 1994 U.S. census data, where the task is to predict whether the annual income of an individual is above \$50,000. We consider the gender of each individual as the sensitive attribute and train the logistic regression as the classification model.
  • The ENEM dataset [enem-do2018instituto] contains about 1.4 million samples from Brazilian college entrance exam scores along with student demographic information. We follow [beyond-alghamdi2022beyond] to quantized the exam score into 2 or 5 classes as label, and consider race as sensitive attribute. As [beyond-alghamdi2022beyond] used a random subset of 50K samples, we instead sample 100K data points to construct our federated dataset. We train multilayer perceptron (MLP) as the classification model.
  • The CelebA dataset [celeba-zhang2020celeba] is a facial image dataset consists of about 200k instances with 40 binary attribute annotations. We identify the binary feature smile as target attributes which aims to predict whether the individuals in the images exhibit a smiling expression. The race of individuals is chosen as sensitive attribute. We train Resnet18 [resnet-he2016deep] on CelebA as the classification model.
The determination of sensitive attributes and labels on three datasets has been verified significant in previous research [beyond-alghamdi2022beyond, benchmark-han2024ffb].
@@ C.1.2 Baselines
We compare the performance of FedFACT with traditional FedAvg [fedavg] and five SOTA methods tailored for calibrating global and local fairness in FL, namely FairFed [fairfed], FedFB [fedfb-zeng2021improving], FCFL [FCFL-cui2021addressing], praFFL [Preference-Aware-federated-group-fairness], and the method in [cost-local-global-fair-duan2025cost], denoted as Cost in our experiments.
@@
  • FedFB presents a FairBatch-based approach [fairbatch] to compute the coefficients of FairBatch parameters on the server. This method integrates global reweighting for each client into the FedAvg framework to fulfill fairness objectives.
@@
  • [cost-local-global-fair-duan2025cost] proposed a convex‐programming-based post-processing framework that characterizes and enforces the minimum accuracy loss required to satisfy specified levels of both local and global fairness constraints in multi-class federated learning by approximating the region under the ROC hypersurface with a simplex and solving a linear program, denoted as Cost in our experiments.
@@ C.1.3 Parameter Settings
For the fairness‐control parameters, e.g., the parameter $\lambda$ in praFFL [Preference-Aware-federated-group-fairness] and the global and local fairness constraints in Cost [cost-local-global-fair-duan2025cost], we impose stringent fairness requirements on the model in our overall comparative experiments, and we adjust the parameters governing the fairness metrics in the Pareto‐curve experiments.
@@ C.1.4 Experiments Compute Resources
C.2 Discussion about FedFACT and LoGoFair [logofair-zhang2025logofair]
LogoFair [logofair-zhang2025logofair] is designed for binary‐classification in federated learning under both global and local fairness constraints, seeking the Bayes‐optimal classifier. By deriving a closed‐form solution for the fair Bayes classifier, LogoFair reformulates the post‐processing fairness adjustment as a bilevel optimization problem jointly solved by the server and clients, which is an approach conceptually analogous to our post‐processing framework. In binary classification, FedFact and LogoFair both target Bayes-optimal classifiers under constraints disparity metrics expressed in linear form. Theoretically, for an identical fairness metric, our Bayes‐optimal fair classifier characterization covers that of LogoFair. Consequently, we refrain from performing a comparative evaluation of the two approaches.
Our method differs by defining the loss at the client level, thereby achieving lower estimation error than the local group-specific objective in [logofair-zhang2025logofair]. Crucially, by formulating the post‐processing model over the probabilistic simplex instead of restricting outputs to the unit interval $[0,1]$ in the binary case, our framework achieves enhanced scalability and naturally adaptable to multi‐group, multiclass settings.
@@ C.3 Heterogeneous Split of Client Distribution
We propose a partitioning method that introduces heterogeneous correlations between the sensitive attribute $A$ and label $Y$, thereby further elucidating the trade-off between global fairness and local fairness [AGLFOP-hamman2024demystifying].
@@
This approach can be regarded as a generalization of the synergy‐level‐based heterogeneous split in [AGLFOP-hamman2024demystifying] to the multi‐client setting, where the AA-$Y$ correlation for each client is governed by a parameter randomly drawn from $a,b]\subseteq[0,1]$, thereby yielding a more pronounced balance between global fairness and local fairness. Throughout the experimental evaluation, we set $\gamma_{k}\in[0.2,\,0.8]$ to guarantee that every client has a sufficient number of sensitive group samples to assess local fairness.
@@ Appendix D Detailed Experiments Results
D.1 Comparison Result and binary EOP criterion
Parato Curves of DP. We have already presented the numerical comparison between our proposed method and the baselines in the main text; here, we report the Pareto curves illustrating the trade‐off between global fairness and accuracy. More precisely, we compare the trade‐off between accuracy and the global fairness measure, as well as the trade‐off between accuracy and the local fairness measure, as a function of the fairness constraint.
The Pareto curve for the global DP criterion is shown in Figure 2 where the horizontal axis denotes accuracy and the vertical axis represents the fairness metric. Consequently, models located closer to the upper‐right corner exhibit superior accuracy-fairness trade‐offs. As illustrated in Figure 2, our method outperforms all existing state-of-the-art approaches when comparing accuracy against either global fairness in isolation.
@@ D.1 Comparison Result and binary EOP criterion
Parato Curves of EOP. In Figure 3, we illustrate the Pareto curve for the Equalized Odds (EO) criterion-accuracy. Because EOP enforces tighter constraints than DP, precise adherence in a federated context requires large per-group sample counts at each client. Hence, we also compare the global EOP here. Our framework still exceeds all state-of-the-art baselines in trading off accuracy against fairness.
@@
D.2 Details for Multi-Class Classification
Multi-Class fair datasets. We illustrate how FedFACT performs on multi-class prediction using CelebA and ENEM. For CelebA, with ’Gender’ still serving as the sensitive attribute, We employ the binary attributes “Smile” and “Big_Nose” to construct a multiclass task by mapping their joint values $\{0,1\}\times\{0,1\}$ onto a four‐class label set $\{0,1,2,3\}$, thereby formulating a multiclass classification problem on the CelebA dataset. These attributes are commonly used in centralized machine learning literature [posthoc, celeba-attribute-pmlr-v202-zhu23n] to construct fairness-aware classification tasks. For ENEM, we follow [beyond-alghamdi2022beyond] to quantize the Humanities exam score to 5 classes. In order to guarantee adequate per-group sample sizes at each client in heterogeneous settings for fairness evaluation (or some clients only hold less than 10 samples for specific group under heterogeneous partitioning), we adopt the four race labels “Branca,” “Preta,” “Parda,” and “Amarela” from the Race attribute as the sensitive groups. These datasets are partitioned into five clients under a heterogeneous split with $\gamma=1$.
Evaluation. In terms of baselines, only the Cost [cost-local-global-fair-duan2025cost] algorithm is theoretically applicable to fairness optimization in multiclass federated learning scenarios. However, their experiments and code are limited to binary classification, and have already been used as binary baselines for comparison with our method. Consequently, we focus exclusively on reporting FedFACT’s performance along with FedAvg in multiclass fairness, establishing it as a pioneering approach in this setting.
@@ D.5 Efficiency and Scalability Study
Efficiency. We evaluate the communication efficiency of FedFACT by monitoring its performance across varying numbers of communication rounds $T$. As illustrated in Figure 4, the post‐processing method, built upon a fully trained pre‐trained model, consistently achieves convergence in fewer than 10 communication rounds, underscoring its high efficiency. The in‐processing method likewise converges in under 40 iterations; given that it requires training the federated model from scratch, this performance is comparable to the convergence speed of FedAvg, making it highly effective compared to existing federated learning algorithms.
Figure 4: Communication Effectiveness Analysis. The convergence rates of both the in‐processing (top row) and post‐processing (bottom row) methods with respect to communication rounds on Compas, Adult, CelebA, and ENEM datasets.
@@
Overall, whether employing the in‐processing or post‐processing method, all three performance metrics rapidly converge to stable values across each of the four datasets, empirically confirming both the communication efficiency and the overall effectiveness of FedFACT.
@@ Appendix E Broader Impacts and Limitations
Broader Impacts. This paper addresses critical fairness issues in FL. By embedding fairness constraints at both the global and client levels, our framework delivers models that distribute accuracy more equitably, bolstering user confidence and mitigating bias amplification. The contributions of this research enhance user satisfaction and promote social equity. This fairness‐aware approach extends readily to high‐stakes classification tasks beyond FL: for instance, clinical decision support in hospital networks, vision‐based detection systems, and financial fraud alerts. Integrating fairness into decentralized model training promotes privacy‐preserving, equitable AI, helps satisfy emerging regulatory requirements, and encourages broader adoption of responsible machine learning across diverse application domains.
Limitations. The primary limitation of FedFACT is the fairness representation, which contains the linear disparities such as communly used DP, EOP and EOP criteria, but it excludes some nonlinear formulations of fairness, e.g. Predictive Parity [compas-dressel2018accuracy] and individual fairness [causal-fairness-NEURIPS2024_4ae5cdac]. Moreover, based on our generalization‐error analysis, although the proposed method enables a controllable accuracy-fairness trade‐off for a given fairness metric, it still requires a sufficiently large local sample size to accurately estimate local fairness (whereas global fairness demands only an adequate overall sample size). While our empirical results compare favorably against existing approaches, exploiting dataset characteristics to optimize fairness may reduce the sample complexity needed for local fairness optimization. Addressing these limitations remains an important avenue for future work.